(20, 240) and (15,450) lies on line


1
Expert's answer
2021-08-02T10:53:22-0400

The points "(20, 240)" and "(15,450)" lies on the line given by the following equation:


"\frac{x-15}{20-15}=\frac{y-450}{240-450}," which is equivalent to the following equations


"\frac{x-15}{5}=\frac{y-450}{-210}"


"y-450=-42(x-15)"


"y=-42x+1080"